A salt dissolves in water by absorbing 1.5 kJ of heat energy from the surrounding 52.0 g of solution. If the solution is initially at 24.0°C, what is the final temperature of the solution?           

      Csoln = 4.18 J/g·deg

Please show steps
Answer #1

Qabsorbed = 1.5 kj

= m*S*DT

DT = T2-T1    (T2>T1)

1.5*10^3 = 52*4.18*(24-T1)

T1 = final temperature of the solution = 17.1 C